OpenStudy (anonymous):

1)Find the indicated limit, if it exists. lim x-> -1 f(x), f(x)= { x+1 x≤ -1, 1-x x≥-1} -1 2 The limit does not exist. 0 2.)Find the indicated limit, if it exists. { 7-x^2 x<0, 7 x=0 10x+7 x>0} The limit does not exist. 7 -10 -17

OpenStudy (anonymous):

@Miss.SweetiePie @Australopithecus

OpenStudy (anonymous):

@uri

OpenStudy (anonymous):

@amistre64 help

OpenStudy (michele_laino):

hint: the graph of your function is: 

OpenStudy (michele_laino):



OpenStudy (anonymous):

i really dont know :(((( @Michele_Laino

OpenStudy (michele_laino):

we have these limits: \[\Large \mathop {\lim }\limits_{x \to - {1^ - }} f\left( x \right) = 0,\quad \mathop {\lim }\limits_{x \to - {1^ + }} f\left( x \right) = 2\]

OpenStudy (michele_laino):

so we have a jump type discontinuity